Superior Markets, Inc., operates three stores in a large metropolitan area. A segmented absorption costing income statement for the company for the last quarter is given below:

Superior Markets, Inc.
Income Statement
For the Quarter Ended September 30
Total North
Store
South
Store
East
Store
Sales $ 4,600,000 $ 920,000 $ 1,840,000 $ 1,840,000
Cost of goods sold 2,530,000 565,000 953,000 1,012,000
Gross margin 2,070,000 355,000 887,000 828,000
Selling and administrative expenses:
Selling expenses 849,000 247,400 323,000 278,600
Administrative expenses 463,000 122,000 174,900 166,100
Total expenses 1,312,000 369,400 497,900 444,700
Net operating income (loss) $ 758,000 $ (14,400 ) $ 389,100 $ 383,300

The North Store has consistently shown losses over the past two years. For this reason, management is giving consideration to closing the store. The company has asked you to make a recommendation as to whether the store should be closed or kept open. The following additional information is available for your use:

  1. The breakdown of the selling and administrative expenses that are shown above is as follows:

Total North
Store
South
Store
East
Store
Selling expenses:
Sales salaries $ 230,400 $ 62,200 $ 68,600 $ 99,600
Direct advertising 181,000 67,000 88,000 26,000
General advertising* 69,000 13,800 27,600 27,600
Store rent 305,000 85,000 118,000 102,000
Depreciation of store fixtures 24,000 6,200 7,600 10,200
Delivery salaries 25,800 8,600 8,600 8,600
Depreciation of delivery
equipment
13,800 4,600 4,600 4,600
Total selling expenses $ 849,000 $ 247,400 $ 323,000 $ 278,600

*Allocated on the basis of sales dollars.

Total North
Store
South
Store
East
Store
Administrative expenses:
Store managers' salaries $ 94,000 $ 29,000 $ 38,000 $ 27,000
General office salaries* 69,000 13,800 27,600 27,600
Insurance on fixtures and inventory 41,000 12,300 17,000 11,700
Utilities 81,120 26,860 24,880 29,380
Employment taxes 62,880 17,040 21,420 24,420
General office—other* 115,000 23,000 46,000 46,000
Total administrative expenses $ 463,000 $ 122,000 $ 174,900 $ 166,100

*Allocated on the basis of sales dollars.

  1. The lease on the building housing the North Store can be broken with no penalty.

  2. The fixtures being used in the North Store would be transferred to the other two stores if the North Store were closed.

  3. The general manager of the North Store would be retained and transferred to another position in the company if the North Store were closed. She would be filling a position that would otherwise be filled by hiring a new employee at a salary of $12,800 per quarter. The general manager of the North Store would continue to earn her normal salary of $13,800 per quarter. All other managers and employees in the North store would be discharged.

  4. The company has one delivery crew that serves all three stores. One delivery person could be discharged if the North Store were closed. This person’s salary is $5,600 per quarter. The delivery equipment would be distributed to the other stores. The equipment does not wear out through use, but does eventually become obsolete.

  5. The company pays employment taxes equal to 15% of their employees' salaries.

  6. One-third of the insurance in the North Store is on the store’s fixtures.

  7. The “General office salaries” and “General office—other” relate to the overall management of Superior Markets, Inc. If the North Store were closed, one person in the general office could be discharged because of the decrease in overall workload. This person’s compensation is $6,900 per quarter.

Required:

1. How much employee salaries will the company avoid if it closes the North Store?

2. How much employment taxes will the company avoid if it closes the North Store?

3. What is the financial advantage (disadvantage) of closing the North Store?

4. Assuming that the North Store's floor space can’t be subleased, would you recommend closing the North Store?

5. Assume that the North Store's floor space can’t be subleased. However, let's introduce three more assumptions. First, assume that if the North Store were closed, one-fourth of its sales would transfer to the East Store, due to strong customer loyalty to Superior Markets. Second, assume that the East Store has enough capacity to handle the increased sales that would arise from closing the North Store. Third, assume that the increased sales in the East Store would yield the same gross margin as a percentage of sales as present sales in the East store. Given these new assumptions, what is the financial advantage (disadvantage) of closing the North Store?
Answer #1
1 Salaries which can be avoided,
Sales salaries 62200
Delivery salaries 5600
Store management salaries (29000-13800) 15200
Salary of new manager 12800
General office compensation 6900
Total 102700
2 Employment taxes which can be avoided=Salaries which can be avoided*15%=102700*15%=$ 15405
3 $ $
Gross margin lost if north store is closed -355000
Savings in cost:
Store rent 85000
Sales salaries 62200
Direct advertising 67000
Delivery salaries 5600
Store management salaries 15200
Salary of new manager 12800
General office compensation 6900
Insurance (12300*2/3) 8200
Utilities 26860
Employment taxes 15405 305165
Decrease in net profit if North store is closed -49835
Financial disadvantage of closing North store=$ 49835
4 $ $
Gross margin lost if north store is closed -355000
Savings in cost:
Store rent 0
Sales salaries 62200
Direct advertising 67000
Delivery salaries 5600
Store management salaries 15200
Salary of new manager 12800
General office compensation 6900
Insurance (12300*2/3) 8200
Utilities 26860
Employment taxes 15405 220165
Decrease in net profit if North store is closed -134835
Financial disadvantage of closing North store=$ 134835
North store should not be closed since it results in a financial disadvantage of $ 134835
5 Gross margin earned from East store
Gross margin % of east store=Gross margin/Sales=828000/1840000=0.45=45%
Sales transferred from North to East store=1/4 th of sales=920000*1/4=$ 230000
Gross margin earned from sales transferred=230000*45%=$ 103500
If floor space can't be subleased,Then threre is no savings in store rent.
Savings in cost=$ 220165
Gross margin lost if north store is closed -355000
Gross margin earned from East store (Note:2) 103500
Net gross margin lost -251500
Savings in cost 220165
Net disadvantage of closing north store -31335